### v1

@@ Abstract
To address these gaps, we propose TabulaTime, a novel multimodal deep learning framework that integrates clinical risk factors with air pollution data to enhance ACS risk prediction. TabulaTime introduces three major innovations: (1) Multimodal feature integration to combine time-series air pollution data with clinical tabular data for improved predictive accuracy; (2) PatchRWKV, a new module for automatic extraction of complex temporal patterns, overcoming limitations of traditional feature engineering and maintaining linear computational complexity; and (3) Enhanced interpretability using attention mechanisms to reveal interactions between clinical and environmental factors.
Our experimental evaluation demonstrated that TabulaTime achieves a 20.5% improvement in accuracy compared to the best-performing traditional machine learning model (CatBoost) and surpasses other models, including Random Forest and LightGBM, by 20.5% to 32.2%. The integration of air pollution data improves accuracy by 10.1%, emphasising the importance of environmental factors. The PatchRWKV module showed advanced performance in time-series analysis, outperforming the state-of-the-art methods such as Multi-Layer Perceptrons (MLPs) based methods (LightTS and Dlinera), Convolutional Neural Networks (CNNs) based methods (TimesNet), Recurrent Neural Networks (RNNs) based methods (LSTM, GRU), and Transformers based methods (Autoformer, PatchTST) across various tasks while maintaining efficiency. Additionally, feature importance analysis identified key predictors of ACS, such as previous angina, systolic blood pressure, and pollutants like PM10 and NO2, highlighting the critical role of environmental-clinical interactions in ACS risk.
@@ 1 Introduction
    Multimodal feature integration and ACS prediction: We propose a novel TabulaTime deep learning framework for multimodal feature integration and ACS prediction, combining extracted time-series features from air pollution datasets with ACS risk factors obtained from clinical tabular data. This integration enables the model to conditionally learn attention maps based on the multimodal features, thereby enhancing both predictive performance and interpretability.
@@ 2.3 Environmental Factors and ACS
CNNs are traditionally used for image processing but have been adapted for time series analysis[44]. The key advantage of CNNs is their ability to capture local patterns and hierarchical features through convolutional operations.
@@ 3.2 Input Embedding
In this work, the input embedding consists of two parts: embedding clinical tabular data and embedding air pollution time series data.
Embedding clinical tabular data involves several crucial steps as follows: First, handling missing values is essential; we fill missing data with the mean value of the respective attribute to minimise the impact on model prediction. Next, we encode categorical variables using One-Hot encoding to convert them into binary vectors, allowing the model to process categorical information effectively. Following this, data normalisation and standardisation are applied, transforming features to have a mean of 0 and a standard deviation of 1, which improves model performance by ensuring consistency in data scale. Finally, all features are concatenated into a single feature vector, serving as the input for model integration, enabling the deep learning framework to utilise the comprehensive dataset efficiently.
Embedding air pollution time series data includes patching and patch embedding, which will be described in detail in a later section.
@@ 3.3.1 Patching
The patching process consists of two stages: normalisation and patching. Instance normalisation has recently been shown to mitigate the distribution shift effect between training and testing data. Therefore, we first normalise each univariate time-series pollution dataset to have a zero mean and unit standard deviation before applying patching. Each input univariate time series is then tokenised through patching, with patches being either overlapping or non-overlapping, depending on the patch length (P) and stride (S). These patches are projected into a lower-dimensional space using a learnable projection matrix, reducing the number of input tokens and computational complexity. In this study, we use hourly air pollution data and aim to capture localised features of pollution for each day. Consequently, we set the patch size to 24 and do not use overlapping patches.
@@ 3.3.2 Patch Embedding
In patch embedding, each patch is embedded into a higher-dimensional vector by applying a linear transformation to flatten it. Then each patch is represented as a vector, which captures the essential information of that patch that can be processed by the RWKV model.
@@ 3.3.3 RWKV (Receptance Weighted Key Value) Encoder
Where h is the number of heads. Finally, an output gating mechanism controls the flow of information from the recurrent unit to the next layer or the final output. This gating is implemented using the SiLU activation function and receptance. The output vector $o_{t}$ per head is given by:
$o_{t}=\left(\operatorname{SiL}\mathrm{U}\left(g_{t}\right)\odot\text{ LayerNorm }\left(r_{t}\cdot wkv_{t}\right)\right)W_{o}$ (4)
Where LayerNorm operates on each of h heads separately. This gating ensures that only relevant information is passed forward, improving the model’s ability to make accurate predictions by filtering out noise and focusing on significant data.
@@
$\begin{gathered}k_{t}^{\prime}=W_{g}^{\prime}\cdot\left(\mu_{k}^{\prime}\odot x_{t}+\left(1-\mu_{k}^{\prime}\right)\odot x_{t-1}\right)\\ r_{t}^{\prime}=W_{r}^{\prime}\cdot\left(\mu_{r}^{\prime}\odot x_{t}+\left(1-\mu_{r}^{\prime}\right)\odot x_{t-1}\right)\\ v_{t}^{\prime}=\operatorname{ReLU}^{2}\left(k_{t}^{\prime}\right)\cdot W_{v}^{\prime}\\ o_{t}^{\prime}=\operatorname{Sigmoid}\left(r_{t}^{\prime}\right)\odot v_{t}^{\prime}\end{gathered}$ (5)
@@ 3.4 Multimodal Feature Integration and Prediction
Feature integration and prediction in the TabulaTime model combine extracted features from time-series air pollution data with tabular clinical data for ACS risk prediction(Figure 4). This model integrates tabular data embedding feature representations of time-series data into an attention module, enabling the network to learn attention maps conditionally based on the tabular data. This integration enhances the network’s ability to pinpoint what, where, and when to focus on in the data, leading to improved performance in tasks involving both time series and tabular data. Specifically, the tabular data is embedded into the same dimension as the feature representations of the time-series data and passed through shared layers. This allows the attention maps to be computed conditionally on the tabular data, facilitating a more nuanced and effective feature integration process.
Give the input $X\in\mathbb{R}^{C}$, the feature integration can be formulated as:
$\hat{X}_{c}=\sigma\left(W_{2}\cdot\delta\left(W_{1}\cdot X_{c}\right)\right)_{c}\cdot X_{c}$ (6)
Where $W_{1}\in\mathbb{R}^{\frac{c}{r}\times c}$ and $W_{2}\in\mathbb{R}^{C\times\frac{c}{r}}$ are the weight matrices of the two fully connected layers. $\sigma$ denotes the ReLU activation function. $\sigma\left(W_{2}\cdot\delta\left(W_{1}\cdot X_{c}\right)\right)_{c}$ refers to the generated attention map. As a final step, a MLP is added at the end of the model to predict the ACS risk.
@@ 4.1 Experimental Design
    This experiment evaluates the predictive accuracy of TabulaTime for Acute Coronary Syndrome (ACS) risk. We test the model using varying air pollution data durations (3, 7, 10, and 14 days) to identify the optimal time window. Subsequently, we compare its performance against traditional machine learning models, including Random Forest, LightGBM [48], and CatBoost[49], under two scenarios: (a) incorporating air pollution data, and (b) excluding air pollution data. These comparisons quantify the significance of environmental factors in enhancing predictive accuracy. Additionally, the generalizability of TabulaTime are assessed across varying time periods and air conditions.
@@
    This study aim to understand the interplay between clinical and environmental factors in ACS STEMI VS NSTEMI prediction. By analyzing feature importance scores, we identified the key predictors and their relative contributions, enhancing the model’s interpretability and offering valuable insights for targeted interventions. The feature importance is calculated using permutation feature importance[55]. This method assesses the increase in the model’s prediction error when a feature’s values are shuffled. If shuffling a feature increases the model error, the feature is considered important, as the model relies on it for predictions. Conversely, if shuffling does not affect the model error, the feature is deemed unimportant, as the model ignores it for predictions. We first calculate feature importance as the drop in the model’s validation metric when a feature value is randomly shuffled. Then, we calculate step importance to show the significance of each time period.
@@ 4.2 Dataset Description
The study utilises two primary datasets to evaluate the performance of the TabulaTime model: 1) Salford MINAP dataset which contains clinical data from the Myocardial Ischaemia National Audit Project (MINAP) collected in Salford, UK. 2) Air Pollution Dataset which contains hourly measurements of various air pollutants collected from the Salford Eccles monitoring station in Salford, Greater Manchester, United Kingdom. 3) Additionally, to evaluate the performance of PatchRWKV, we have also used three public available time series datasets including Weather [56], Heartbeat Monitoring Dataset and Self-Regulation of Slow Cortical Potentials (SCP) Dataset.
@@ 4.2.1 Salford MINAP Dataset
In this work, 21 risk factors including patient demographics, clinical history, medications, vital signs and diagnostic indicators are selected to predict the ACS risk and differentiate between STEMI and NSTEMI based on their relevance and significance in clinical and environmental contexts.
Factors such as previous Acute Myocardial Infarction (AMI), angina, hypertension, hypercholesterolemia, peripheral vascular disease, cerebrovascular disease, asthma/COPD, chronic renal failure, heart failure, smoking status, diabetes, previous PCI (Percutaneous Coronary Intervention), and previous CABG (Coronary Artery Bypass Graft) are included. These factors are crucial as they provide a comprehensive view of a patient’s medical history and risk factors associated with cardiovascular diseases[57, 58]. Use of ACEI or ARB (Angiotensin-Converting Enzyme Inhibitors or Angiotensin II Receptor Blockers), systolic blood pressure (BP), height, weight, BMI (Body Mass Index), family history of coronary heart disease (CHD), creatinine levels, and statin use are considered. These factors help in assessing the ongoing treatment, physiological conditions, and genetic predisposition of patients. The detailed risk factors are shown in Table 1.
Based upon electrical heart tracings – electrocardiograms or ECGs recorded during a heart attack, patients are diagnosed as having suffered either ST-elevation myocardial infarction (STEMI) or non-ST elevation myocardial infarction (NSTEMI). In the annual NCAP report, the terms ‘higher-risk’ and ‘lower-risk’ have been used to differentiate STEMI from NSTEMI.
Medical Conditions and Measurements Counting and calculation formulae Previous AMI (Acute Myocardial Infarction) Yes 381/No 597 Previous Angina Yes 470/No 511 Hypertension Yes 591/No 389 Hypercholesterolaemia Yes 623/No 358 Peripheral Vascular Disease Yes 103/No 877 Cerebrovascular Disease Yes 125/No 855 Asthma/COPD Yes 274/No 704 Chronic Renal Failure Yes 143/No 837 Heart Failure Yes 187/No 792 Smoking Status Ex 384/Current smoker 251/Never 343 Diabetes Yes 298/No 683 Previous PCI (Percutaneous Coronary Intervention) Yes 184/No 797 Previous CABG (Coronary Artery Bypass Graft) Yes 99/No 880 ACEI or ARB (Angiotensin-Converting Enzyme Inhibitors or Angiotensin II Receptor Blockers) Yes 471/No 507 Systolic BP (Blood Pressure) 137.5 ± 27.1 Height 166.6 ± 10.3 Weight 78.1 ± 20.57 BMI BMI= Weight ( $\mathrm{BMI}=\frac{\text{ Weight }}{(\text{ Height })^{2}}$ Family History of CHD (coronary heart disease) Yes 176 /No 785 Creatinine 112.4 ± 90.3 Statin Yes 592/No 386 Age 71.1 ± 13.7 STEMI 123/587
Table 1: Feature used in MINAP dataset.
4.2.2 Air Pollution Dataset

### v2

@@ Abstract
To address these gaps, we propose TabulaTime, a novel multimodal deep learning framework that integrates clinical risk factors with air pollution and climate data to enhance ACS risk prediction. TabulaTime introduces three major innovations: (1) Multimodal feature integration to combine time-series environmental data (air pollution and climate) with clinical tabular data for improved predictive accuracy; (2) PatchRWKV, a new module for automatic extraction of complex temporal patterns, overcoming limitations of traditional feature engineering and maintaining linear computational complexity; and (3) Enhanced interpretability using attention mechanisms to reveal interactions between clinical and environmental factors.
Our experimental evaluation demonstrated that TabulaTime achieves a 20.5% improvement in accuracy compared to the best-performing traditional machine learning model (CatBoost) and surpasses other models, including Random Forest and LightGBM, by 20.5% to 32.2% in differentiating ACS presentation (STEMI vs. NSTEMI). The integration of air pollution and climate data improves accuracy by 10.1%, emphasising the importance of environmental factors. The PatchRWKV module showed advanced performance in time-series analysis, outperforming state-of-the-art methods such as Multi-Layer Perceptrons (MLPs) based methods (LightTS and DLinear), Convolutional Neural Networks (CNNs) based methods (TimesNet), Recurrent Neural Networks (RNNs) based methods (LSTM, GRU), and Transformers based methods (Autoformer, PatchTST) across various classification and forecasting tasks while maintaining linear computational efficiency. Additionally, feature importance analysis identified key predictors, including clinical factors like Systolic Blood Pressure, Symptom-to-Admission Time, and BMI, alongside environmental factors such as PM10 (maximum and average), average NO, and temperature, highlighting the critical role of environmental-clinical interactions in ACS presentation.
@@ 1 Introduction
    Multimodal feature integration and ACS prediction: We propose a novel TabulaTime deep learning framework for multimodal feature integration and ACS prediction, combining extracted time-series features from environmental datasets with ACS risk factors obtained from clinical tabular data. This integration enables the model to conditionally learn attention maps based on the multimodal features, thereby enhancing both predictive performance and interpretability.
@@ 2.3 Environmental Factors and ACS
CNNS Convolutional Neural Networks (CNNs), commonly used for image processing, had also been successfully adapted for time-series analysis, demonstrating strength in capturing local patterns and hierarchical features through convolutional operations [44]. However, standard CNN architectures often struggled with modelling long-range temporal dependencies due to their limited receptive fields. To address this, several recent approaches in the Long-Term Time Series Forecasting (LTSF) domain introduced architectural innovations that extended temporal modelling capabilities while maintaining computational efficiency. For instance, SCINet [45] employed a recursive downsample–convolve–interact mechanism to extract multi-resolution temporal features. MICN [46] used a multi-scale architecture that combined downsampled convolutions for local features with isometric convolutions for global context, achieving strong performance with linear complexity. TimesNet [47], which we included as a CNN-based baseline in our evaluation, transformed 1D time series into 2D tensors and applied 2D convolutions to capture periodic and structured temporal variations. Despite these advancements, CNNs remained inherently limited in capturing extended temporal relationships—especially in clinical and environmental health contexts where delayed effects, such as those from air pollution exposure prior to ACS onset, were crucial. Extending CNNs’ temporal range typically required deeper or dilated architectures, which increased model complexity and computational cost. These limitations motivated the development of more expressive and scalable frameworks capable of natively modelling long-range dependencies.
@@ 3.2 Input Embedding
In this work, the input embedding consists of two parts: embedding clinical tabular data and embedding air pollution time series data. The specific dimensions and processes are detailed in Section 4.4.2 (Implementation Details).
Embedding clinical tabular data involves several preprocessing steps. First, missing values (approximately 2% in our dataset) were handled using K-Nearest Neighbours (KNN) imputation [51] (Details regarding the selection of KNN and the sensitivity analysis are provided in Section 4.4.2). Next, categorical variables were converted into a numerical format using One-Hot encoding. Subsequently, all numerical features were standardised to have a mean of 0 and a standard deviation of 1 to ensure consistency in scale. Finally, the fully processed feature vector for each patient is embedded into a dense vector representation $X_{tab}\in\mathbb{R}^{D}$ using a multi-layer perceptron (MLP). This embedded vector serves as the input for the multimodal integration stage.
Embedding air pollution time series data involves segmenting the normalised time series into patches and then projecting these patches into an embedding space, resulting in tokens $X_{ts}\in\mathbb{R}^{N\times T\times C}$. This process is described in detail in the following sections (3.3.1 and 3.3.2).
@@ 3.3.1 Patching
The patching process consists of two stages: normalisation and patching. Instance normalisation is applied first to each univariate time-series pollution dataset to mitigate distribution shift effects. Each normalised time series is then tokenised into patches. In this study, using hourly data, we set the patch size P and stride S based on capturing daily patterns (details in Section 4.4.2). These patches are treated as input tokens by the subsequent RWKV encoder.
@@ 3.3.2 Patch Embedding
Following normalisation and patching, each patch (a vector of dimension P) is linearly projected into a higher-dimensional embedding space (dimension C). These embedded patch representations form the time-series input tokens $X_{ts}$ for the PatchRWKV module, structured as $X_{ts}\in\mathbb{R}^{N\times T\times C}$, where N is the number of pollution indicators, T is the number of patches (days), and C is the latent feature dimension. The specific dimensions used are detailed in Section 4.3.2.
@@ 3.3.3 RWKV (Receptance Weighted Key Value) Encoder
Where h is the number of heads. Finally, an output gating mechanism controls the flow of information from the recurrent unit to the next layer or the final output. The output gating is implemented using the sigmoid ($\delta$) of $r_{t}$. The output vector $o_{t}$ per head is given by:
$o_{t}=\left(\delta(r_{t})\cdot wkv_{t}\right)W_{o}$ (4)
where $W_{o}$ represents the weight matrix for the output projection following the gating and attention calculation.
@@
$\begin{gathered}k_{t}^{\prime}=W_{g}^{\prime}\cdot\left(\mu_{k}^{\prime}\odot x_{t}+\left(1-\mu_{k}^{\prime}\right)\odot x_{t-1}\right)\\ r_{t}^{\prime}=W_{r}^{\prime}\cdot\left(\mu_{r}^{\prime}\odot x_{t}+\left(1-\mu_{r}^{\prime}\right)\odot x_{t-1}\right)\\ v_{t}^{\prime}=\operatorname{ReLU}^{2}\left(k_{t}^{\prime}\right)\cdot W_{v}^{\prime}\\ o_{t}^{\prime}=\operatorname{\delta}\left(r_{t}^{\prime}\right)\odot v_{t}^{\prime}\end{gathered}$ (5)
@@ 3.4 Multimodal Feature Integration and Prediction
Feature integration and prediction in the TabulaTime model combine the embedded clinical tabular data vector with the sequence of extracted features from the time-series air pollution data for ACS risk prediction(Figure 4). This model integrates tabular data embedding feature representations of time-series data into an attention module, enabling the network to learn attention maps conditionally based on the tabular data. This integration enhances the network’s ability to pinpoint what, where, and when to focus on in the data, leading to improved performance in tasks involving both time series and tabular data.
Let $X_{tab}\in\mathbb{R}^{D}$ be the embedded vector representation of the clinical tabular data, processed as described in Section 3.2 and projected to the dimension D. $X_{ts}\in\mathbb{R}^{N\times T\times C}$ represent the sequence of feature vectors output by the PatchRWKV module (where N is the number of air pollution types, T is the effective sequence length (number of days), and C is the feature dimension). This tensor is first flattened across the N and T dimensions, and then global average pooling is applied across the flattened dimension, resulting in a 1D vector $X_{ts\_pooled}\in\mathbb{R}^{N\times T}$.
The model integrates these representations using an attention mechanism. Give the input $X\in\mathbb{R}^{D+N\times T}$, the feature integration can be formulated as:
$\hat{X}=\sigma\left(W_{2}\cdot\delta\left(W_{1}\cdot X\right)\right)\cdot X$ (6)
Where $W_{1}$ and $W_{2}$ are the weight matrices of the two fully connected layers. $\delta$ denotes the ReLU activation function.$\sigma$ denotes the sigmoid activation function. $\sigma\left(W_{2}\cdot\delta\left(W_{1}\cdot X_{c}\right)\right)$ refers to the generated attention map. As a final step, a MLP is added at the end of the model to predict the ACS risk.
@@ 4.1 Experimental Design
    This experiment evaluates the predictive accuracy of TabulaTime for Acute Coronary Syndrome (ACS) risk. We test the model using varying air pollution and climate data durations (3, 7, 10, and 14 days) to identify the optimal time window. Subsequently, we compare its performance against traditional machine learning models, including Random Forest, LightGBM [52], and CatBoost[53], under two scenarios: (a) incorporating environmental data, and (b) excluding environmental data. These comparisons quantify the significance of environmental factors in enhancing predictive accuracy. Additionally, the generalisability of TabulaTime are assessed across varying time periods and air conditions.
@@
    This study aims to evaluate the predictive performance of a novel deep learning framework (TabulaTime) that integrates selected clinical risk factors and time-series environmental data for differentiating ACS presentation (STEMI vs. NSTEMI). In this experiment, we assess the model’s ability to identify significant features and their contributions to the prediction of ACS risk. By analysing feature importance scores, we identified the key predictors and their relative contributions, enhancing the model’s interpretability and offering valuable insights for targeted interventions. The feature importance is calculated using permutation feature importance[58]. This method assesses the increase in the model’s prediction error when a feature’s values are shuffled. If shuffling a feature increases the model error, the feature is considered important, as the model relies on it for predictions. Conversely, if shuffling does not affect the model error, the feature is deemed unimportant, as the model ignores it for predictions. We first calculate feature importance as the drop in the model’s validation metric when a feature value is randomly shuffled. Then, we calculate step importance to show the significance of each time period.
@@ 4.2 Dataset Description
The study utilises three primary datasets to evaluate the performance of the TabulaTime model: 1) Salford MINAP dataset, 2) Air Pollution and Climate Dataset and 3) Additionally, to evaluate the performance of PatchRWKV, we have also used three public available time series datasets including Weather [59], Heartbeat Monitoring Dataset and Self-Regulation of Slow Cortical Potentials (SCP) Dataset.
@@ 4.2.1 Salford MINAP Dataset
In this work, 21 risk factors including patient demographics, clinical history, medications, vital signs and diagnostic indicators are selected to predict the ACS risk and differentiate between STEMI and NSTEMI based on their relevance and significance in clinical and environmental contexts. These 21 selected risk factors represent a combination of data types pertinent to ACS risk and presentation:
Baseline and Historical Data: Including patient demographics (age, sex, BMI derived from height and weight), documented pre-existing conditions and medical history (e.g., Previous AMI, Previous Angina, Hypertension, Hypercholesterolaemia, Diabetes, Chronic Renal Failure, Heart Failure, Cerebrovascular Disease, Peripheral Vascular Disease, Asthma/COPD, Family History of CHD), prior interventions (Previous PCI, Previous CABG), relevant lifestyle factors (Smoking Status), and chronic medication usage (ACEI or ARB, Statin). These factors reflect the patient’s condition prior to the specific ACS event which are crucial as they provide a comprehensive view of a patient’s medical history and risk factors associated with cardiovascular diseases[60, 61].
Admission Data: Including key physiological measurements typically obtained upon presentation or during the initial assessment for the ACS event, such as Systolic Blood Pressure and serum Creatinine levels. Additionally, the time from symptom onset to therapy was calculated by determining the duration between the recorded ’Onset of Symptoms’ time and the ’Admission Date/Time’ for each patient.
This comprehensive variable set allows the model to learn from both long-term antecedent risk factors and indicators present at the time of the acute clinical presentation, in conjunction with the preceding environmental data.The detailed risk factors are shown in Table 1.
Based upon electrical heart tracings – electrocardiograms or ECGs recorded during a heart attack, patients are diagnosed as having suffered either ST-elevation myocardial infarction (STEMI) or non-ST elevation myocardial infarction (NSTEMI).
Medical Conditions and Measurements Counting and calculation formulae Previous AMI (Acute Myocardial Infarction) Yes 381/No 597 Previous Angina Yes 470/No 511 Hypertension Yes 591/No 389 Hypercholesterolaemia Yes 623/No 358 Peripheral Vascular Disease Yes 103/No 877 Cerebrovascular Disease Yes 125/No 855 Asthma/COPD Yes 274/No 704 Chronic Renal Failure Yes 143/No 837 Heart Failure Yes 187/No 792 Smoking Status Ex 384/Current smoker 251/Never 343 Diabetes Yes 298/No 683 Previous PCI (Percutaneous Coronary Intervention) Yes 184/No 797 Previous CABG (Coronary Artery Bypass Graft) Yes 99/No 880 ACEI or ARB (Angiotensin-Converting Enzyme Inhibitors or Angiotensin II Receptor Blockers) Yes 471/No 507 Systolic BP (Blood Pressure) 137.5 ± 27.1 Height 166.6 ± 10.3 Weight 78.1 ± 20.57 BMI BMI= Weight ( $\mathrm{BMI}=\frac{\text{ Weight }}{(\text{ Height })^{2}}$ Family History of CHD (coronary heart disease) Yes 176 /No 785 Creatinine 112.4 ± 90.3 Statin Yes 592/No 386 Age 71.1 ± 13.7 SymptomToAdmissionTime -8673 to 15131 Hours STEMI 123/587
Table 1: Features used in MINAP dataset.
4.2.2 Air Pollution and Climate Dataset
@@
In addition to air pollution, hourly climate data, specifically temperature (measured in degrees Celsius), was incorporated into the analysis. This data was obtained from the Meteostat (https://meteostat.net/en/) public data source, using readings from a weather station geographically close to the study area (Manchester Airport station, identifier 03330). The climate data covers the same temporal range as the air pollution datasets to ensure alignment for the multimodal analysis. Temperature data was included as it is a known environmental factor that can influence cardiovascular events.
Crucially, both the air pollution and climate datasets were temporally aligned with the clinical data from the Salford MINAP dataset. For each patient’s admission date, the corresponding preceding environmental data (e.g., for the 10 days prior) was extracted and matched, ensuring that the model analyses the environmental conditions leading up to the ACS event.